OpenStudy (anonymous):

Medals and fan! quick help please! calculate |-3 + 4i|, thats all! i just need it to be explained

OpenStudy (anonymous):

@Compassionate you were recommended, can you help me?

jimthompson5910 (jim_thompson5910):

|-3 + 4i| is basically asking for the modulus of -3+4i

jimthompson5910 (jim_thompson5910):



OpenStudy (bradely):

if a+bi is complex number then |a+bi| =sqrt(a^2+b^2) |-3+4i|=sqrt((-3)^2+(4i)^2) =sqrt(9+16i^2) Here i^2 =-1 =sqrt(9-16) =sqrt(-7) =sqrt(7i^2) =isqrt(7) Source: http://www.mathskey.com/question2answer/